Operations Report – November 11th

Crew 182 Operations Report  11Nov2017

SOL: 8

Name of person filing report:  Carmen Atauconcha

Non-nominal systems:

WATER HEATER: It is not working.  In the morning, we had to take a shower with cold water. The water heater is not heating the water.  We opened the entire water pipe, but it still did not heat.

COMMUNICATION RADIOS: The constantly noise of one of the radios stops.
SECONDARY AIRLOCK: The door of the secondary airlock is gotten out-off-square.  I am saying that it is wonky since we arrived on Mars.
ATV’s: The blue ATV is not working. When we try to turn it on, it does not answer.

Generator—  It worked the whole night since 6pm of 10Nov2017 until 8:00 am of 11Nov2017.  Turn on at 5.15 pm 11Nov2017.  Operand.




9:11 am 100%
10:20 am 100%
11:20 am 100%
12:20 am 100%
13:20 am 100%
14:20 pm 100%
15:50 pm 98%
16:20 pm 97%
16:45 pm 94%
17:04 pm 91%


Diesel – 55 %

Propane – 50 psi

Gasoline (5 Gallon containers for ATV) – 1/4 gallons. In addition the tank of gasoline inside the ATV’s is filled.

Water (trailer) –  0 gallons

Water (static) – 258  gallons

Trailer to Static Pump used –   no

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  36.6 gallons

Toilet tank emptied:  yes

ATV’s Used: No

Oil Checked:  No

Atv # Fuel Used Gals:  0  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 0 hours

Deimos used:  No

Hours: 0

Beginning charge:  –

Ending charge:  –

Currently charging: Yes

HabCar used? No

Notes and Comments:
Today we had an advance in the topic about recycling water. The number of gallons of water that we used today is less than the gallons of water used yesterday.  And, I know that tomorrow is going to be better.
Summary of  Hab operations:

We receive a new digester for the toilet.  And we wait that it removes the bad smell of the toilet. We will keep you post.

The area of the tools in the Hab was a mess. Today we clean and organize this place.  Also, in the afternoon, we flushed the toilet with the new digester.   From another hand, we are going to receive the communication radios on Monday.
Finally, today we received some food.  This is great for us. Thanks¡

Summary of GreenHab operations:

We turned on the heater at 18:12 pm with a temperature of 47 °F.  At the same time, we are cleaning this area to make it beautiful.

Summary of ScienceDome operations:

We received a new shop vac. It has been installed in the laboratory. Also, there were cables going through the laboratory that we order to keep the security in the laboratory.

Summary of health and safety issues:

The HOS officer needs two fire alarms and lamps for the science dome and for the greenhouse. In the morning we fixed the stair of the secondary airlock.  Now, that stair is more stable.

Questions, concerns and requests to Mission Support:

We need you to repair the heater water, since it still does not work.

For the next week at the most on Monday, we are going to need water.

I have a doubt, which is the minimum limit for the propane tank and the diesel tank?

Where does the used water go?

When are we going to receive water?

Copyright © The Mars Society. All rights reserved. | Main Site